OpenStudy (anonymous):

Unique Uniforms Company charges $17.75 per cap plus a design fee of $30. logos Direct Inc charges $15.50 per cap plus a delivery fee of $48. For what number of caps ordered would the cost be the same from both vendors?

OpenStudy (anonymous):

Let n be the number of caps. Solve the following for n:\[30+\frac{1775}{100}n=\frac{1550}{100}n+48 \]\[n=8 \]

OpenStudy (anonymous):

Did you just randoly divide by 100 to get the answer?

OpenStudy (anonymous):

Multiply each side by 100 to get rid of the fractions. By using fractions in the equation setup, the solution will always be symbolically exact.

OpenStudy (anonymous):

Okay so you added 30 plus 17.75 then divided that by 100 then you added 15.50 plus 48 and subtracted the two answers?

OpenStudy (anonymous):

\[100\left(30+\frac{1775}{100}n\right)=\left(\frac{1550}{100}n+48\right)100 \]\[3000+1775 n=4800+1550 n\]\[1775 n-1550 n=4800-3000 \]\[225 n=1800\]\[n=\frac{1800}{225}=8 \]

OpenStudy (anonymous):

Ooooh okay I get it now thanks soo much :) you really helped me out

OpenStudy (anonymous):

You are welcome.

OpenStudy (anonymous):

And thank you for the medal.

OpenStudy (anonymous):

No problem you deserved it :)